OpenStudy (anonymous):

x-2y=6 solve for y a. y=x-3 b. y=-x+3 c. y=1/2x-3 d. y=-1/2x+3

OpenStudy (hoblos):

c

OpenStudy (anonymous):

c. y=1/2x-3

OpenStudy (anonymous):

I get y=(x-6)/2

OpenStudy (anonymous):

same as y=1/2x-3

OpenStudy (anonymous):

Oh.....You mean (1/2)*x-3 I read it as 1/(2x-3)